8 is 2 and 1/2 of what number?
2/3 divide by 5?

Found 2 solutions by ptaylor, Alan3354:
Answer by ptaylor(2198)

8 is 2 and 1/2 of what number?
Let x=the number
2 1/2 is equal to 5/2, so:
(5/2)x=8 multiply each side by 2
5x=16 divide each side by 5
x=3 1/5 = 16/5-------------the number
CK
(16/5)(5/2)=8
8=8
2/3 divide by 5?
5/(2/3)=x
First multiply the numerator and denominator of the left side by (3/2). This will get rid of the complex fraction.
5*(3/2)/(2/3)*(3/2)=x
15/2/1=x
x=15/2=7 1/2
CK
(15/2)*(2/3)=5
5=5
Hope this helps---ptaylor

Answer by Alan3354(69443)

8 is 2 and 1/2 of what number?
I don't know what that means.
---------------
2/3 divide by 5?
(2/3) / (5/1)
Invert the divisor and multiply
= (2/3) * (1/5) = 2/15